A student is performing the experiment of resonance Column. The diameter of the column tube is 4 cm. The frequency of the tuning fork is 512 Hz. The air temperature is 38 °C in which the speed of sound is 336 m/s. The zero of the meter scale coincides with the top end of the Resonance Column tube. When the first resonance occurs, the reading of the water level in the column is ?

Grade:11

1 Answers

Vikas TU
3 years ago
V/4(l+e)=f 
⇒l+e=V/4f
⇒l=(V/4f)−e
Here e=(0.6)r=(0.6)(2)=1.2cm
So l=336×10^2/4×512−1.2=15.2cm........................................